Circular No. 7126

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1999 E1 Weidong Li and M. Modjaz, Department of Astronomy, University of

California at Berkeley, report the discovery of a comet in the course of the Lick Observatory Supernova Search (cf. IAUC 6627; with the participation of J. Y. King, M. Papenkova, R. A. Johnson, A. Friedman, R. R. Treffers and A. V. Filippenko). The object was found automatically by the Katzman Automatic Imaging Telescope (KAIT), recorded by Modjaz as a supernova candidate, and recognized by Li as a comet. Li then used the equipment to make confirmatory observations three nights later. The comet shows an apparent tail in p.a. ∼ 120^{o}. Available unfiltered CCD astrometry:

1999 UT  α_{2000}  δ_{2000}  m_{1}
Mar. 13.17569  6^{h}24^{m}42.^{s}94  +4^{o}41′24′′.2  16.5
16.14473  6 25 40.92  +5 16 38.0  16.8
16.16436  6 25 41.29  +5 16 51.7  16.7
16.19431  6 25 41.93  +5 17 12.7  16.8

4U 0115+63 W. A. Heindl and W. Coburn, University of California, San Diego,

report: “The current outburst of the 3.6-s x-ray pulsar 4U 0115+63 (IAUC 7116) has been monitored with daily RXTE pointings since Mar. 3. The flux (2–10 keV) has increased from 5.6 × 10^{−9} (Mar. 3) to 1.1 × 10^{−8} erg cm^{−2} s^{−1} on Mar. 12. Fitting the Mar. 3 joint PCA/HEXTE spectrum reveals cyclotron absorption features at ∼ 14 and 21 keV. Because the phase-averaged spectra are complex due to pulse-phase spectral variability, these line energies are not well-determined. A cyclotron absorption feature was first seen in this source with HEAO 1 (Wheaton et al. 1979, Nature 282, 240). In addition to the short monitoring observations, we have made two long pointings on Mar. 7.16–7.45 and 11.89–12.32 UT. The second of these spanned periastron passage (according to the ephemeris of Bildsten et al. 1997, Ap.J. Suppl. 113, 367) at Mar. 11.95. During this observation, the source exhibited large oscillations with a timescale of ∼ 500 s. The oscillation rms deviations were ∼ 30 percent of the mean flux. Such slow oscillations are not consistent with the beat-frequency QPO model whose expected frequencies are of order 5 Hz. The oscillations may have been present during 2500 s of the Mar. 7 observation, but at greatly reduced amplitude. Additional observations are encouraged. The RXTE/ASM real-time light

curve is available at http://heasarc.gsfc.nasa.gov/xte weather/.”

1999 March 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green